9 - x square over y square over 3 - x over y
Answer by oscargut(2103)

[(9-x^2)/y^2]/[(3-x)/y]=
[(9-x^2)/y^2]=(3-x)(3+x)y^2/(3-x)y=(3+x)y
Answer: (x+3)y